In a recently administered IQ test, the scores were distributed normally, with mean 100 and standard deviation 15. What proportion of the test takers scored between 70 and 130?

A. About 68%;

B. About 84%

C. About 95%

D. About 99.5%
Answer #1

According to 68-95-99.7 rule, 68% of data lies between 1 standard deviation from mean, 95% between two standard deviation from mean & about 99.7% lies between 3 standard deviation from mean.

Q. H:100, 6515 , 70= 100-2x15=4-26, 130= 4+26 - P(70<x<130): PU-26 < X < 4+20) X: Test score : © About 95% [68-95-99.7 Rule]

If you found solution satisfactory, please rate or do comment in comment section for review or further clarification. Thank you